OpenStudy (anonymous):

sin of 15 degrees... ?

OpenStudy (anonymous):

I assume u mean without a calculator. One way would be sin(a+b)=sin(a)cos(b)+cos(a)sin(b). Set a = 45 and b = -30 sin(15 degrees) = sin(45)cos(30) - cos(45)sin(30)